# Adaptive Intelligence: leveraging insights from adaptive behavior in animals to build flexible AI systems

### v2

@@
Our world is built on a series of predictions made in our mind. That is, to us, the world is a construct, created conceptually from a series of millions of predictions harbored in the neural code. When we walk on the beach, we accumulate photons with our eyes, sound waves with our ears, we feel the wind on our face, and the sand shifting under our feet, and all of this comes together to give us the perception of a warm, sunny day with ocean waves crashing on the shore. However, our brain sits in the dark confines of our skull and tens of billions of neurons communicate through electrical activity in a bath of chemicals. As incoming sensory information is processed in the brain at different times, the brain imposes a lag in order to compile it, thrusting our perception into the past—tens of precious milliseconds behind reality. Nevertheless, we need to act quickly, and our brains have evidently evolved solutions to overcome these delays.
@@
Moreover, we need to adapt to environments marked by unpredictability, continuous change, and infrequent repetition of decision-making scenarios. The uncertainty, variability, and complexity of these settings demand adaptive strategies, crucial for learning about local changes – such as the shifting sand under our feet. Dating back decades, researchers have developed new frameworks to understand (decompile) how this global prediction system may work (1, 2, 3, 4). A core tenant has been that in order to overcome sensory-processing delays we must build internal models of the world. We use said models to then make predictions about the sensory and motor consequences of our actions (5, 6, 7, 8). But it remains largely unknown how the brain enables adaptive behavior, and how we can take inspiration from the brain to build more adaptive artificial systems (9, 10).
Dating at least as far back as Aristotle, there has been a philosophical and scientific question of how our brains build models of the world (11). How is it that we can combine sensory information into perception? How long does this take? Are we always tens, or even hundreds of milliseconds behind reality? How could our brains make predictions about the world in order to act faster? Why do we mostly notice when the world violates our predictions? – if the ocean were suddenly quiet, we would instantly worry we had lost our hearing and not doubt that the ocean had completely changed. While this is familiar in everyday life, we need to game-ify this approach to study it in the laboratory.
What is the neural basis of this behavior? This historically has been difficult to answer, because we are always under-sampling the data and often want repeated measures. The mouse brain has an estimated 70 million neurons, yet even the most state-of-the-art technology allows for recording up to 1 million individual neurons at a time (12) (and on the order of only 3Hz resolution), while most labs are on the order of hundreds to thousands with two-photon microscopy or high-density arrays such as Neuropixels (13, 14). Thus, if one aims to study how a neuron encodes a given stimulus, you can repeatedly subject the animal to the stimulus and build encoding and decoding models (15) to gain insight into what the neural activity represents. However, if you want to understand how the brain adapts, there is an inherent problem with this approach. As Heraclitus infamously quipped, “No man ever steps in the same river twice, for it’s not the same river and he’s not the same man”; the neural code is changing dynamically so averaging across trials becomes, at best, an over simplification that would miss core-principles of neural computations.
Thus, to study the neural mechanisms underlying adaptive behavior we need three ingredients: (1) behavioral paradigms with measurable rapid-learning (within-session); (2) large-scale neural recordings; and (3) data analysis methods that consider time-series and can leverage many neurons, either within a single subject or across subjects, for single trial, continuous analysis. The first I will cover here, while the second I will not cover but rather refer the reader to reviews covering the rapid pace of technology development in two-photon imaging, flexible neural implants, and high-density recordings (16, 17, 18). For the third, I will briefly discuss algorithms for time-series analysis, but then venture further into how AI and neuroscience can both mutually enable studying and building adaptive systems.
@@
Figure 1: Rapid Learning in animals: from few-shot to updating of internal model-based learning. (a) Adapted from Rosenberg et al. 19: Four sample bouts from one mouse (B3) into the maze at various times during the experiment (time markings at bottom). The trajectory of the animal’s nose is shown; time is encoded by the color of the trace. The entrance from the home cage and the water port are indicated in the top left panel. (b) Adapted from El-Gaby et al. 20: Task design: animals learned to navigate between 4 sequential goals on a 3×3 spatial grid-maze. Reward locations changed across tasks but the abstract structure, 4 rewards arranged in an ABCD loop, remained the same. Bottom left: When allowed to learn across multiple sessions, animals readily reached near-optimal performance in the last 20 trials, as demonstrated by comparing path length between goals to the shortest possible path. Bottom right: Performance improved across the initial 20 trials of each new task. This improvement was markedly more rapid for the last 5 tasks compared to the first 5 tasks. (c) Adapted from (7): Top Diagram of the joystick adaptation task structure: 75 baseline trials, 100 perturbation trials, and 75 “washout” (i.e., null field) trials. Bottom Left: an example session showing the average baseline trajectory (black), the first perturbation trial (sea green), average of first ten (green), and the last ten perturbation trials (purple), the first ten washout trials (orange) and the first washout trial (red). The numbers indicate trials. Right: average trajectories for all sessions (n = 27 sessions, from n = 7 mice) with temporally aligned averaging. Dashed sea green line is the average first perturbation, and the dashed red line is the average of the first five washout trials. Solid lines are average paw path ± SEM in the direction of the perpendicular deviation, with the same color scheme as the left panel; solid green is the first 25 trials during the perturbation epoch, solid purple is the last 25 trials of the perturbation epoch, and solid red line is the average of the first 25 washout trials.
Adaptive animal & neural behavior
@@
There has been a concurrent push for more naturalistic behaviors, more complex behaviors, and tasks that enable within-session learning. Aside from new large-scale neural recording technology, a part of this is due to the advances in markerless tracking of behavior (reviewed in Mathis and Mathis (21)), which has made digitizing movements much more tractable. What behaviors are being developed? While I cannot cover all progress in the field, I want to highlight several new approaches. Freely moving paradigms for rodents have rapidly advanced to encompass complex stimuli such as moving (threatening) novel objects (22), real-time VR worlds (23), labyrinth mazes (19), extended multi-area home cages (24, 25), and 3D environments such as those used to study depth estimation (26). In several settings they have also built trial-based assays that smartly constrain these naturalistic settings in order to be able to directly probe learning across epochs.
@@
An elegant thread of work across multiple groups aims to build tasks for zero- or few-shot learning in complex spatial worlds. For example, Rosenberg et al. (19) developed a task where mice must seek rewards in a labyrinth (Figure 1a). They found that mice can not only make up to 2000 decisions per hour, but after unsupervised exploration of the maze they could few-shot learn to find the water spout (in around 10 tries) (19). Notably, this is a learning rate that is 1000-fold higher than 2AFC tasks that are typically deployed in rodent studies of decision-making. Another example is showing that mice can zero-shot learn new sequences of rewards (such as go to location A, B, D, then C) (20). Mice were trained on multiple tasks with a shared underlying structure that organized a sequence of goals, though the specific goal locations varied. Strikingly, the mice learned this common structure, allowing them to make zero-shot inferences on the first trial of new tasks (Figure 1b). Intriguingly, they also found neurons in medial prefrontal cortex that acted as task-structured memory buffers, namely, they tracked the progression along behaviorally-relevant steps (20).
@@
In sensorimotor control there has been a long line of work on rapid learning. So-called motor adaptation studies have been developed where visual or proprioceptive information is perturbed such that within-session, an animal must adapt. This has not only spurred the development of new neural analysis tools (27, 28, 29, 15), but a host of behavioral tasks. For example, visuomotor rotations have been used in humans and non-human primate studies to study how they can account for sensorimotor discrepancies in only a few hundred trials (30, 31, 32, 27, 33). The same principle is used for changing environmental dynamics of a manipulandum that causes deviations in a limb-movement tasks (34, 35, 7, 36). Notably, these tasks are specifically designed to measure both adaptive learning, and the formation of internal models. There is always a period of baseline control movements, perturbations, and a return to baseline-condition that allows researchers to behaviorally measure if an internal model was updated (Figure 1c).
@@
Neurally, evidence shows that motor and sensory areas can change their tuning properties during the course of learning (37, 38, 36, 39, 40). A study by Sun et al. (38) showed that during a motor adaptation task in macaques, motor cortex (M1) can effectively index memories of the hand-force required for learning to adapt. Notably, the neural subspace most predictive of hand forces changed during the period before movement (preparatory), specifically during the learning epoch. In a neural dimension orthogonal to this force-predictive subspace, they identified a uniform shift across all movement directions, including those unaffected by learning. Intriguingly this uniform shift remained after exposure to the force field, reflecting an updated internal model. Other works also show new evidence of state-changes that persist across learning (36, 8), and temporally-resolved prediction errors (41, 36). Influential work on visual cortex (V1) has shown the existence of ‘mismatch’ neurons when the expected visual feedback is disrupted (41), and similar prediction errors have been found in somatosensory (S1), motor (M1), and frontal areas of cortex (40, 36).
@@
Some of the best evidence for causally showing how neurons adapt has come from brain-machine-interface (BMI) studies that require the subject to directly alter neural firing in order to change something in the external world (like a cursor on a screen). Closed-loop systems, like calcium-based BMIs (caBMIs) or with electrical stimulation, have been used to study the timescale and subsets of neurons that can be used, as not all cell types have been found to be equally adaptable. Fundamental work has revealed that even a few neurons can be leveraged for decoding, and for BMI-guided feedback (27, 42). Notably, they often live in discrete subspaces of the neural dynamics. Vendrell-Llopis et al. (43) pushed this further to link these subspaces to cell-types. They trained mice to modulate the activity of either intratelencephalic (IT) or pyramidal tract (PT) neurons for reward. They found that mice learned to control PT neuron activity more quickly and effectively than IT neuron activity. This intriguingly could be related to the anatomical connectivity and differing inputs.
@@
Related to cell types, the anatomy of the brain is important to consider, as this too could have direct implications for the future design of adaptive AI systems. Several areas in the brain have this incredible structure, – layers in the cerebral (neo-) and cerebellar cortex – which could be a critical part of new architectures for AI systems. There is a long-standing debate about representation emergence from data vs. architecture constraints (49, 50), but it is increasingly clear that architecture is critically important and linked to function.
@@
Fundamental work dating back to the 1970’s is worth revisiting. Vernon Mountcastle describes a framework of the organizational principle of the neocortex (51). He delineates evidence that the major functional division in the neocortex is not whether an area is “sensory” or “motor,” but rather the vertical neocortical column constitutes the basic computational unit, and the input-output pattern merely dictates the space of information it acts upon—namely, the “auditory”, “visual” or “motor neocortex” has the same cellular scaffold and should only considered a particular region based on the type of sensory input. Nonetheless, this architectural bias gives rise to computations on information with brain area-specific information.
@@
Form ties to function, and one emerging hot topic is how cortical circuits implement learning from prediction errors across cell types and layers. Building on the initial discovery of sensory prediction errors in cortex (41), additional works have begun to record from subtypes of neurons such as parvalbumin (PV), VIP, SST, and excitatory neurons in order to form a more complete model of how both excitatory and inhibitory neurons across layers could implement the learning rule (45, 52, 44) (Figure 2a). This work in cortex follows earlier work that discovered reward prediction errors (RPEs) in the midbrain (within the ventral tegmental area) (53, 48, 47), and they are also part of an inhibitory-gated circuit, where GABAergic neurons play a critical role in the computation of RPEs (54) (Figure 2b).
@@
Recent work aims to unify a view for how a hierarchical implementation of prediction errors (including RPEs) could aid in learning across sensory, cognitive, and motor systems (55). Tsai et al. (55) et al. found that during a cue-guided reward task, layer 2/3 somatosensory neurons showed enhanced responses to reward-predictive stimuli, and this led to reduced reward-prediction errors and increased confidence in predictions. Following rule reversal, the lateral orbitofrontal cortex, via VIP interneurons, signaled context-prediction errors, reflecting a loss of confidence. The work suggests a hierarchical interaction of prediction errors across cortical regions, with top-down signals modulating sensory cortex activity Notably, prior work finds limited roles for only “pure” reward-based learning in motor learning, but it can be combined with sensory prediction errors to shape performance (32, 7, 56). I envision that future work will more deeply test how various learning signals are used concurrently across systems. For example, tasks that have different types of prediction errors can be developed that might conflict, forcing the subject to weight different errors in order to guide future actions.
In the context of learning and links to AI, another critical view on cell types and anatomy is neural development. For example, cells in the cortex migrate to their final resting location in different times. Moreover, winner-takes-all synapses form, such as at the neuro-muscular-junction of alpha motor neurons and muscles. There is also a complex and highly regulated developmental program in the spinal cord of transcription factors and hox genes – the genetics lay out the global (rostral-caudal) and even local (dorsal-ventral zones with a level) connections. Additionally, there is the so-called genomic bottleneck to consider, which aids in determining early innate abilities of animals – from the immediate suckling reflex to the slightly more time-intensive but still impressive ability of a horse to stand and walk a few hours after birth (57).
@@
Moreover, minimally in cortex, early traveling spontaneous waves form the basis of early connections between cortical neurons. Intriguingly, new work has highlighted how the transformer architecture naturally gives rise to waves, which are a critical feature in the brain (58). Waves are not only seen in early development (as mentioned), but also in steady-state – where they are thought to keep the brain’s subregions interconnected and sharing neural dynamics that could encode global states and arousal levels.
@@
New tools – i.e., the third ingredient – are critical for the study of adaptive behavior. Computational neuroscience has made large advances in the past decade, fueled by the need to better understand ensembles of neurons (17, 59, 60, 61) and by leveraging the concurrent innovations in deep learning, which is reviewed in Hurwitz et al. (62) and Mathis et al. (15). In brief, algorithms such as variational autoencoders, transformers, contrastive learning-optimized neural networks, and diffusion models are already making big strides in our ability to combine data across subjects, measure model multi-region interactions, and ‘decode the brain’ (15).
@@
For review, I want to briefly highlight several tools that enable time-series analysis in single neurons and neural populations, as these are critical for adaptive learning studies, and I suspect will become more popular tools for studying AI systems. One of the most impactful tools in the last 20 years has been the adoption of generalized linear models (GLMs) for neuroscientific applications (63, 64). Here, a model is trained using a feature matrix (can be one or many features) with the goal of predicting the spikes of neurons. While GLMs can come in many flavors, as an example, to handle variability in behavior, models are trained and tested on different splits of the data (also for cross validation). Predictions from these test sets are combined, and a pseudo-R2 is calculated. This does require that behavior is repeated, but also allows for flexibility; i.e., there is no direct trial-averaging or behavior-triggered averaging of the neural code.
@@
For ensembles of neurons, several methods have been used, ranging from PCA for measuring neural trajectories (65, 66), contrastive learning paired with consistency metrics to measure changes in the latent neural dynamics across learning (36, 29), to nonlinear dynamical systems for task- and task-irrelevant metrics (67, 68). These are more deeply covered in reviews elsewhere (62, 15), but I will note that several approaches can be used for single-trial dynamics and have been especially impactful in motor learning and adaptation studies in non-human primates (61, 69). In particular, work by Azabou et al. (70) introduce a training framework and architecture designed to model population dynamics in large-scale neural recordings, which tokenizes individual spikes to capture fine temporal structures and constructs a latent representation of neural population activity. They trained this on nearly 100 hours of data and show excellent cross-task generalizing in primate reaching studies. One limitation of this approach is the requirement of labels at test-time, but in practice this is nearly always available.

### v3

@@ Introduction
Our world is built on a series of predictions made in our mind. That is, to us, the world is a construct, created conceptually from a series of millions of predictions harbored in the neural code. When we walk on the beach, we accumulate photons with our eyes, sound waves with our ears, we feel the wind on our face, and the sand shifting under our feet, and all of this comes together to give us the perception of a warm, sunny day with ocean waves crashing on the shore. However, our brain sits in the dark confines of our skull and tens of billions of neurons communicate through electrical activity in a bath of chemicals. As incoming sensory information is processed in the brain at different times, the brain imposes a lag in order to compile it, thrusting our perception into the past – tens of precious milliseconds behind reality. Nevertheless, we need to act quickly, and our brains have evidently evolved solutions to overcome these delays.
@@
Moreover, we need to adapt to environments marked by unpredictability, continuous change, and infrequent repetition of decision-making scenarios. The uncertainty, variability, and complexity of these settings demand adaptive strategies, crucial for learning about local changes – such as the shifting sand under our feet. Dating back decades, researchers have developed new frameworks to understand (decompile) how this global prediction system may work (1, 2, 3, 4). A core tenet has been that in order to overcome sensory-processing delays we must build internal models of the world (also called world models). We use the models to make predictions about the sensory and motor consequences of our actions (5, 6, 7, 8). But it remains largely unknown how the brain enables adaptive behavior, and how we can take inspiration from the brain to build more adaptive artificial intelligence (AI) is of growing interest (9, 10, 11). Specifically, AI here broadly refers to systems or machines that aim to mimic biological intelligence by performing tasks that typically require cognition, such as problem solving, decision making, learning, and language processing. Technically, it encompasses machine learning (ML), deep learning, natural language processing (NLP), and computer vision. In this Perspective, I review progress in studying biological intelligence and use this to propose how we can enhance AI-based agentic systems.
@@
Figure 1: Rapid Learning in animals: from few-shot to updating of internal model-based learning. (a) Adapted from Rosenberg et al. 12: Four sample bouts from one mouse (B3) into the maze at various times during the experiment (time markings at bottom). The trajectory of the animal’s nose is shown; time is encoded by the color of the trace. The entrance from the home cage and the water port are indicated in the top left panel. (b) Adapted from El-Gaby et al. 13: Task design: animals learned to navigate between 4 sequential goals on a 3×3 spatial grid-maze. Reward locations changed across tasks but the abstract structure, 4 rewards arranged in an ABCD loop, remained the same. Bottom left: When allowed to learn across multiple sessions, animals readily reached near-optimal performance in the last 20 trials, as demonstrated by comparing path length between goals to the shortest possible path. Bottom right: Performance improved across the initial 20 trials of each new task. This improvement was markedly more rapid for the last 5 tasks compared to the first 5 tasks. (c) Adapted from (7): Top Diagram of the joystick adaptation task structure: 75 baseline trials, 100 perturbation trials, and 75 “washout” (i.e., null field) trials. Bottom Left: an example session showing the average baseline trajectory (black), the first perturbation trial (sea green), average of first ten (green), and the last ten perturbation trials (purple), the first ten washout trials (orange) and the first washout trial (red). The numbers indicate trials. Right: average trajectories for all sessions (n = 27 sessions, from n = 7 mice) with temporally aligned averaging. Dashed sea green line is the average first perturbation, and the dashed red line is the average of the first five washout trials. Solid lines are average paw path ± SEM in the direction of the perpendicular deviation, with the same color scheme as the left panel; solid green is the first 25 trials during the perturbation epoch, solid purple is the last 25 trials of the perturbation epoch, and solid red line is the average of the first 25 washout trials.
Internal models for adaptive behavior
Dating at least as far back as Aristotle, there has been a philosophical and scientific question of how our brains build models of the world (14). How can we combine sensory information into perception? How long does this take? Are we always tens or even hundreds of milliseconds behind reality? How could our brains make predictions about the world in order to act faster? Why do we mostly notice when the world violates our predictions? – if the ocean were suddenly quiet, we would immediately worry that we had lost our hearing and not doubt that the ocean had completely changed. While this is familiar in everyday life, we need to turn this experience into a controllable task to study it in the laboratory.
What is the neural basis of this adaptive behavior? This historically has been difficult to answer because we are always under-sampling the neural data that underlies the behavior and often want repeated measures. The mouse brain has an estimated 70 million neurons, yet even the most state-of-the-art technology allows for recording up to 1 million individual neurons at a time (15) (and on the order of only 3Hz resolution), while most labs can record in the order of hundreds to thousands of neurons with two-photon microscopy or high-density arrays such as Neuropixels (16, 17). Thus, classically, to study how a neuron encodes a given stimulus the animal is repeatedly subjected to the stimulus and then we build encoding and decoding models (18) to gain insight into what the neural activity represents.
However, if you want to understand how the brain learns, there is an inherent problem with this repeated-trial approach. As Heraclitus infamously quipped, “No man ever steps in the same river twice, for it’s not the same river and he’s not the same man”. Beyond the clear changes that can occur across trials during learning, the neural code can change even in steady-state given that the output behavior is rarely truly identical. Thus, averaging across trials likely misses key principles of neural computation. Since behavior is never exactly repeated anyhow, investigating learning — where behavior changes in predictable and measurable ways due to task design — may offer crucial insights into the neural basis of adaptive behavior. Namely, we can control the environmental factors that drive learning and examine the neural code across time. This not only lends itself to more ecologically-relevant tasks (the animals must be able to learn), but it’s becoming a tractable approach given modern technological advances in large-scale neural recording and joint neural and behavioral analysis (19, 20, 21, 18).
Adaptive animal and neural behavior
@@
There has been a concurrent push for experimental designs that incorporates more naturalistic behaviors, more complex behaviors, and tasks that enable within-session learning. Aside from new large-scale neural recording technology, a part of this is due to the advances in markerless tracking of behavior (reviewed in Mathis and Mathis (22)), which has made digitizing movements much more tractable. What behavioral paradigms are being developed? While I cannot cover all progress in the field, I want to highlight several new approaches. Freely moving paradigms for rodents have rapidly advanced to encompass complex stimuli such as moving (threatening) novel objects (23), real-time VR worlds (24), labyrinth mazes (12), extended multi-area home cages (25, 26), and 3D environments such as those used to study depth estimation (27). In several settings they have also built trial-based assays that smartly constrain these naturalistic settings in order to be able to directly probe learning across epochs.
@@
An elegant thread of work across multiple groups aims to build tasks for zero- or few-shot learning in complex spatial worlds. Zero-shot learning is a term adapted from ML that describes the ability of a model to generalize to new tasks or concepts without any specific training examples (28). Few-shot means that it can learn from only a few examples. For example, Rosenberg et al. (12) developed a task where mice must seek rewards in a labyrinth (Figure 1a). They found that mice can not only make up to 2000 decisions per hour, but after unsupervised exploration of the maze they could few-shot learn to find the water spout (in around 10 tries) (12). Notably, this is a learning rate that is 1000-fold higher than two-alternative forced choice (2AFC) tasks that are typically deployed in rodent studies of decision-making, meaning this task allows one to better study rapid learning. Another example is showing that mice can zero-shot learn new sequences of rewards (such as go to location A, B, D, then C) (13). Mice were trained on multiple tasks with a shared underlying structure that organized a sequence of goals, though the specific goal locations varied. Strikingly, the mice learned this common structure, allowing them to make zero-shot inferences on the first trial of new tasks (Figure 1b). Intriguingly, they also found neurons in medial prefrontal cortex that acted as task-structured memory buffers, namely, they tracked the progression along behaviorally-relevant steps (13). Such schemata could perhaps be mapped to memory-replay in neural networks (see below).
Animal intelligence is, of course, not limited to laboratory animals. Dating back to early comparative cognition studies (29) there has been a large body of work studying intelligence in a myriad of animals. Crows, for example, exhibit tool use and causal reasoning, with New Caledonian crows even manufacturing tools to retrieve food (30, 31). Bees demonstrate numeracy and complex communication through the waggle dance, encoding spatial information about food sources (32) and can solve “puzzles" such as pulling a string to move a food source closer (33). Various domesticated animals also exhibit human-aligned intelligence. Horses show social learning and can interpret human emotional cues (34). Dogs display theory of mind and word comprehension that rival that of young children (35). Such findings even sparked a reoccurring virtual animal AI Olympics challenge (36).
@@
In sensorimotor control there has been a long line of work on rapid learning, which is called motor adaptation. For clarity, here I define learning as the acquisition of a skill (and thus learning a new internal model), while adaptation is using a learned internal model and updating it. Motor adaptation studies have been developed where visual or proprioceptive information is perturbed such that within-session, an animal must adapt. This has not only spurred the development of new neural analysis tools (37, 38, 39, 18), but a host of behavioral tasks. For example, visuomotor rotations have been used in humans and non-human primate studies to study how they can account for sensorimotor discrepancies in only a few hundred trials (40, 41, 42, 37, 43). The same principle is used for changing environmental dynamics of a manipulandum that causes deviations in a limb-movement tasks (44, 45, 7, 46). Notably, these tasks are specifically designed to measure both adaptive learning, and the formation of internal models. There is always a period of baseline control movements, perturbations, and a return to baseline-condition that allows researchers to behaviorally measure if an internal model was updated (Figure 1c).
@@
Neurally, evidence shows that motor and sensory areas can change their tuning properties during the course of learning (47, 48, 46, 49, 50). A study by Sun et al. (48) showed that during a motor adaptation (learning) task in macaques, motor cortex (M1) can “index” memories of the hand-force required for the process of learning to adapt in the form of movement readiness potentials that occur prior to movement execution. Notably, the neural subspace most predictive of hand forces changed during the period before movement (preparatory), specifically during the learning epoch. In a neural dimension orthogonal to this force-predictive subspace, they identified a uniform shift across all movement directions, including those unaffected by learning. Intriguingly this uniform shift remained after exposure to the force field, reflecting an updated internal model. In addition, other works also show new evidence of state-changes that persist across learning (46, 8), and temporally-resolved prediction errors (51, 46). Influential work on visual cortex (V1) has shown the existence of ‘mismatch’ neurons when the expected visual feedback is disrupted (51), and similar prediction errors have been found in somatosensory (S1), motor (M1), and frontal areas of cortex (50, 46).
@@
Some of the best evidence for causally showing how neurons adapt has come from brain-machine-interface (BMI) studies that require the subject to directly alter neural firing in order to change something in the external world (like a cursor on a screen). Closed-loop systems, like calcium-based BMIs (caBMIs, BMIs driven by decoded optical activity readout of calcium fluorescence signals in M1) or with electrical stimulation, have been used to study the timescale and subsets of neurons that can be used, as not all neuronal types have been found to be equally adaptable. Fundamental work has revealed that even small numbers of neurons can be leveraged for decoding, and for BMI-guided feedback (37, 52). Notably, they often live in discrete subspaces of the neural dynamics. Vendrell-Llopis et al. (53) pushed this further to link these subspaces to cell-types. They trained mice to modulate the activity of either intratelencephalic (IT) or pyramidal tract (PT) neurons for reward. They found that mice learned to control PT neuron activity more quickly and effectively than IT neuron activity. This intriguingly could be related to the anatomical connectivity and differing inputs.
@@
Anatomical links to computations
@@
The anatomy of the brain is important to consider, as this too could have direct implications for the future design of adaptive AI systems. Several areas in the brain have this incredible structure, – layers in the cerebral (neo-) and cerebellar cortex – which could be a critical part of new architectures for AI systems. There is a long-standing debate about representation emergence from data vs. architecture constraints (59, 60), but it is increasingly clear that architecture is critically important and linked to function. Fundamental work dating back to the 1970’s is worth revisiting. Vernon Mountcastle describes a framework of the organizational principle of the neocortex (61). He delineates evidence that the major functional division in the neocortex is not whether an area is “sensory” or “motor,” but rather the vertical neocortical column constitutes the basic computational unit, and the input-output pattern merely dictates the space of information it acts upon—namely, the “auditory”, “visual” or “motor neocortex” has the same cellular scaffold and should only considered a particular region based on the type of sensory input. Nonetheless, this architectural bias gives rise to computations on information with brain area-specific information.
@@
Form ties to function, and one emerging hot topic is how cortical circuits implement learning from prediction errors across cell types and layers. Building on the initial discovery of sensory prediction errors in cortex (51), subsequent works have begun to record from subtypes of neurons such as parvalbumin (PV), vasoactive intestinal peptide (VIP) and somatostatin (SST) classes of interneurons, and excitatory neurons in order to form a more complete model of how both excitatory and inhibitory neurons across layers could implement the learning rule (55, 62, 54) (Figure 2a). This work in cortex follows earlier work that discovered reward prediction errors (RPEs) in the midbrain (within the ventral tegmental area) (63, 58, 57), and found that they are also part of an inhibitory-gated circuit where GABAergic neurons play a crucial role in the computation of RPEs (64) (Figure 2b).
@@
Recent work aims to unify how a hierarchical implementation of prediction errors (including RPEs) could aid in learning across sensory, cognitive, and motor systems. Tsai et al. (65) found that during a cue-guided reward task, layer 2/3 somatosensory neurons showed enhanced responses to reward-predictive stimuli, and this led to reduced reward-prediction errors and increased confidence in predictions. Following rule reversal, the lateral orbitofrontal cortex, via VIP interneurons, signaled context-prediction errors, reflecting a loss of confidence. The work suggests a hierarchical interaction of prediction errors across cortical regions, with top-down signals modulating sensory cortex activity Notably, prior work finds limited roles for reward-only driven motor learning, but it can be combined with sensory prediction errors to shape performance (42, 7, 66). I envision that future work will more deeply test how various learning signals are used concurrently across systems. For example, tasks that have different types of prediction errors can be developed that might conflict, forcing the subject to weight different errors in order to guide future actions.
Neural development and inductive biases
In the context of learning and the links to AI, another critical angle is neural development. There are natural connections to make that could also inspire new inductive biases in the form of architectures or optimization algorithms in AI systems. For example, during development cells in the cortex migrate to their final resting location at different times (67). This process of spatial and temporal organization, I believe, has parallels in AI, particularly in how artificial neural networks undergo initialization and optimization processes to organize their nodes across layers in deep learning models. Moreover, synapses, such as those at the neuro-muscular junction of alpha motor neurons and muscles, are formed after activity-dependent pruning (68). This mirrors the pruning techniques (so called knowledge distillation in ML (69)) used to improve model performance, where unnecessary connections are discarded during training to enhance efficiency.
There is also a complex and highly regulated developmental program in the spinal cord involving transcription factors and HOX genes (70). This genetic regulation lays out global (rostral-caudal) and even local (dorsal-ventral) connections. Similarly, the initial structure of a neural network is determined by a set of weights and biases, which are refined through training to create the connections necessary for learning. Additionally, the so-called genomic bottleneck plays a key role in determining early innate abilities of animals, from the immediate suckling reflex to the more time-intensive but impressive ability of a horse to stand and walk a few hours after birth (71). In AI, this is akin to pre-trained models or transfer learning, where networks leverage previously learned features to rapidly adapt to new tasks with minimal input.
@@
Moreover, in the cortex, early traveling spontaneous waves shape the basis of early connections between cortical neurons. These waves are thought to synchronize neural activity across different regions of the brain. Intriguingly, new work has highlighted how the transformer architecture naturally gives rise to waves, which are a critical feature in the brain (72). This is reminiscent of how attention mechanisms in transformers enable dynamic, wave-like information flow across different parts of the model, helping the system maintain coherence across large datasets. Waves are seen not only in early development but also in steady-state, where they are thought to maintain global connectivity in the brain, encoding states like arousal and attention. This functional organization in the brain has direct analogs in how AI systems, like attention networks, maintain global contextual understanding across their layers.
@@
New tools are critical for the study of adaptive behavior. Computational neuroscience has made large advances in the past decade, fueled by the need to better understand ensembles of neurons (20, 73, 74, 75) and by leveraging the concurrent innovations in deep learning, which is reviewed in Hurwitz et al. (76) and Mathis et al. (18). In brief, algorithms such as variational autoencoders, transformers, contrastive learning-optimized neural networks, and diffusion models are already making big strides in our ability to combine data across subjects, model multi-region interactions, and ‘decode the brain’ (18). I want to briefly highlight several tools that enable time-series analysis in single neurons and neural populations, as these are critical for adaptive learning studies, and I suspect will become more popular tools for studying AI systems.
@@
One of the most impactful tools in the last 20 years has been the adoption of generalized linear models (GLMs) and generalized additive models (GAMs) for neuroscientific applications (77, 78, 79). Here, a model is trained using a feature matrix (can be one or many features) with the task of predicting the spikes of neurons given the set of features. While GLMs can come in many flavors, as an example, generally to handle variability in behavior models are trained and tested on different splits of the data (also for cross validation). Predictions from these test sets are combined, and a pseudo-R2 is calculated. This does require that behavior is repeated, and also allows for flexibility; i.e., there is no direct trial-averaging or behavior-triggered averaging of the neural code.
@@
For ensembles of neurons, several methods have been used or developed, ranging from PCA for measuring neural trajectories (80, 81), contrastive learning paired with consistency metrics to measure changes in the latent neural dynamics across learning (46, 39), to nonlinear dynamical systems for task- and task-irrelevant metrics (82, 83). These are more deeply covered in reviews elsewhere (76, 18), but I will note that several approaches can be used for single-trial dynamics and have been especially impactful in motor learning and adaptation studies in non-human primates (75, 84). Work by Azabou et al. (85) introduce a training framework and architecture designed to model population dynamics in large-scale neural recordings, which tokenizes individual spikes to capture fine temporal structures and constructs a latent representation of neural population activity. They trained this model on nearly 100 hours of data and show excellent cross-task generalizing in primate reaching studies. One limitation of this approach is the requirement of labels at test-time, but in practice this is nearly always available.